A mixture of 30 pounds of candy sells for $1.10 a pound. The mixture consists of chocolates worth $1.50 a pound and chocolates worth 90¢ a pound. How many pounds of the $1.50 chocolate were used to make the mixture?
A)15
B)10
C)12

Answer by Theo(13342)

your equations are
1.5*x + .9*y = 1.10*30
x + y = 30
solve for x in the second equation:
x = 30-y
substitute for x in the first equation:
1.5*(30-y) + .9*y = 1.10*30
simplify by removing parentheses:
1.5*30 - 1.5*y + .9*y = 1.10*30
subtract 1.5*30 from both sides of the equation:
-.6*y = 1.10*30 - 1.5*30
combine like terms:
-.6*y = -.4*30
solve for y:
y = 20
confirm by substituting in original equation:
x = 10
y = 20
1.5*10 + .9*20 = 1.10*30
simplify:
15 + 18 = 33
confirm:
33 = 33

answer is:

10 pounds of the $1.50 chocolate were used.